OpenStudy (anonymous):

Calculus Help!! Evaluate the indefinite integral as an infinite series ∫sinx /3x dx. Find the first five non-zero terms of series representation centered at x=0. f(x)= What is the radius of convergence? i think it is infinity!! Note: Remember to include an arbitrary constant "C".

OpenStudy (anonymous):

i put f(x)=x/3-x^3/54+x^5/1800-x^7/105840+x^9/9797760+... and R=infinity and it is nor correct!! What did i do wrong?

OpenStudy (kira_yamato):

Perhaps you might like to use this: http://algo.inria.fr/esf/function/SI/SI.pdf

OpenStudy (anonymous):

that's what i did.and my question has to divided it by 3x

OpenStudy (kira_yamato):

Divide by 3 after we found that series... But what's the supposed correct answer?

OpenStudy (anonymous):

i am not sure!!

OpenStudy (anonymous):

find the series (or rather write the series) for sine divide by 3x term by term integrate term by term that should do it right?

OpenStudy (anonymous):

sinx=x-x^3/3!+x^5/5!-x^7/7!+x^9/9! sinx/3x=1/3-x^2/18+x^4/360-x^6/15120+x^8/1088640 ∫sinx /3x dx=x/3-x^3/54+x^5/1800-x^7/105840+x^9/9797760

OpenStudy (anonymous):

R=infinity

OpenStudy (anonymous):

I will do it for you when I come back from my morning walk.

OpenStudy (anonymous):

okay,thanks!!

OpenStudy (anonymous):

i put this too but it show incorrect!!so i don't know what i miss!!

OpenStudy (amistre64):

2n+1 are the odds; not n+1

OpenStudy (amistre64):

\[sin(x)=\sum^{inf}_{n=0}\frac{(-1)^nx^{2n+1}}{(2n+1)!}\] \[\frac{sin(x)}{3x}=\sum^{inf}_{n=0}\frac{(-1)^nx^{2n}}{3(2n+1)!}\]

OpenStudy (anonymous):

but isn't it the first five terms are the same? where you divide sinx by 3x sinx=x-x^3/3!+x^5/5!-x^7/7!+x^9/9! sinx/3x=1/3-x^2/18+x^4/360-x^6/15120+x^8/1088640 ∫sinx /3x dx=x/3-x^3/54+x^5/1800-x^7/105840+x^9/9797760

OpenStudy (amistre64):

\[\int \frac{(-1)^nx^{2n}}{3(2n+1)!}\to\ \frac{(-1)^n}{3(2n+1)!}\int x^{2n}=\frac{(-1)^n x^{2n+1}}{3(2n+1)(2n+1)!}\]

OpenStudy (amistre64):

\[\frac{x}{3*1!}-\frac{x^3}{9*3!}+\frac{x^5}{15*5!}-\frac{x^7}{21*7!}+\frac{x^9}{27*9!}\]

OpenStudy (anonymous):

that's what i put too!! and R=infinity!!but it is incorrect!!

OpenStudy (amistre64):

im not sure what R equals yet :) but ive got to go fer now

OpenStudy (anonymous):

okay.it also Note: Remember to include an arbitrary constant "C".

OpenStudy (anonymous):

but when it doesn't let me to put C

OpenStudy (amistre64):

if your wroking with a mathing program; make sure your answering acording to its "format"

OpenStudy (anonymous):

i am sure it is the format but just doesn't say it is correct:(

OpenStudy (anonymous):

i got it now.i have to put C for the first term!!

OpenStudy (anonymous):

Thanks everyone!!